OpenStudy (anonymous):

Find the surface area of the right prism: A rectangular base 2 ft by 6 ft, height 5 ft

OpenStudy (anonymous):

area= bh if i take 2*6 = 12(5) = 60 ft^3 Is this right?

OpenStudy (anonymous):

SA = LA + 2B is the formula

OpenStudy (anonymous):

LA=lateral area?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The surface area of a prism = 2 × area of base + perimeter of base × H

OpenStudy (anonymous):

92 ft^2 is my answer 12 ft^2 is my base area 16 is my perimeter * height (5) = 80 80 + 12 = 92 ft^2

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

See if your answer matches please

OpenStudy (anonymous):

really?

OpenStudy (amistre64):

you got what you need yet?

OpenStudy (anonymous):

so amistre is here...m leaving now:P

OpenStudy (anonymous):

I think so :) I hope so

OpenStudy (amistre64):

you can stay, i just wanna look over your shoulder lol

OpenStudy (anonymous):

ha thanks

OpenStudy (anonymous):

no m serious to go..stayed as she need help, now a better choice is here :)

OpenStudy (amistre64):

better? where?

OpenStudy (anonymous):

amistre 64, the darling ;P

OpenStudy (amistre64):

lol.... she just likes the way i use my keyboard. :)

OpenStudy (anonymous):

dinabina..i guess we r done with this qs?

OpenStudy (anonymous):

lol... you two its just math!! I appreciate everyone's help. :)

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

amistre u r great in maths... no doubt :)

OpenStudy (amistre64):

i got me moments;) thanx

OpenStudy (anonymous):

@dina..ok

OpenStudy (amistre64):

Surface Area: rectangular base 2 ft by 6 ft, height 5 ft Sa = 2tops + 2lefts + 2fronts .... Sa = 2(6*2) + 2(6*5) + 2(5*2) Sa = 24 +60 +20 Sa = 60 + 44 = 104, right?

OpenStudy (anonymous):

eeks... did you see what I did?

OpenStudy (amistre64):

no..... was it embarassing ?

OpenStudy (anonymous):

formula is same as u applied

OpenStudy (anonymous):

there might be some error in computations

OpenStudy (anonymous):

I found the base area... then the perimeter * by height then added my base area to the lateral area

OpenStudy (anonymous):

got 92

OpenStudy (amistre64):

.... that aint gonna work out tho

OpenStudy (anonymous):

where did I got wrong and how did you apply 2(height)? Just wondering.. :)

OpenStudy (amistre64):

surface area of a box is just adding up all the faces... each face consists of 2 parts to compute area

OpenStudy (amistre64):

6.5 5.2 or 6.2

OpenStudy (anonymous):

righteous

OpenStudy (amistre64):

there are 2 faces that match for each set.....

OpenStudy (amistre64):

2(6.2) + 2(6.5) + 2(5.2) is the total of all faces

OpenStudy (anonymous):

and that would be 104 ft squared not cubed right?

OpenStudy (amistre64):

squared is right

OpenStudy (anonymous):

can you look the problem over with the pentagonal prism... check my answer...please

OpenStudy (anonymous):

oh thats neat

OpenStudy (amistre64):

i dunno ..... im kinda busy doing nothing ... and thats gonna take all day to finish lol

OpenStudy (anonymous):

lol